OpenStudy (anonymous):

find the equation of a line that runs through (7,5) and is parallel to a line with a slope o 3/2

OpenStudy (anonymous):

slope of*

OpenStudy (anonymous):

pàralel means same slope. So slope = 3/2 y-5=3/2(x-7)

OpenStudy (anonymous):

 WE HAVE USED SLOPE OF THE PARALLEL LINE BECAUSE THE SLOPES OF THE PARALLEL LINES ARE EQUAL..:)